OpenStudy (miss.rose):

Can you help me with five (5) questions on Mixed Operations Equations? I will medal and fan!

OpenStudy (anonymous):

ill try

OpenStudy (miss.rose):

Okay, great! First Question: What is the solution to the equation? 7x + 6 = –22 A. x = -64/7 B. x = -4 This one? C. x = -16/7 D. x = 4